# Federico Echenique and Farzad Pourbabaee

### v2

@@ 2.3 Exchange Economies
The set of agents is denoted by $I$ and a typical member is referenced by indices $i,j$, and $k$. We assume throughout that $I$ is finite.
An exchange economy is a mapping $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$, where $\mathcal{E}(i)=(\succeq_{i},\omega_{i})$. Each agent $i\in I$ is described by a preference relation $\succeq_{i}$ on $\mathbb{R}_{+}^{d}$, as well as an endowment vector $\omega_{i}\in\mathbb{R}_{+}^{d}$. In an exchange economy, we use $\mathcal{U}_{i}^{(\varepsilon)}$ to denote the upper contour set $\mathcal{U}_{\succeq_{i}}^{(\varepsilon)}$.
Given an exchange economy $\mathcal{E}$, the aggregate endowment is $\omega\coloneqq\sum_{i\in I}\omega_{i}$. [^5] In the textbook Walrasian setting, $\omega(s)$ represents the total available amount of good $s$ in the economy, and in the Arrow-Debreu market it represents the total contingent amount that agents can receive in state $s$. A profile of acts across agents, say $f=\{f_{i}:i\in I\}\in\mathbb{R}_{+}^{d\times I}$, is called an allocation if $\sum_{i\in I}f_{i}=\omega$. The space of all allocations is denoted by $\mathcal{F}_{\omega}$.
@@
In words, an allocation $f$ in $\mathcal{E}$ is $\varepsilon$-Pareto optimal if there is no redistribution of the aggregate endowment $\omega=\sum_{i}f_{i}$ that would be strictly better for all agents, and that would remain strictly better for all agents after a fraction $\varepsilon$ of consumption is “shaved off” in each state of the world. When $\varepsilon=0$, the definition coincides with the usual notion of weak Pareto optimality. Further properties of the set of $\varepsilon$-Pareto optimal allocations are discussed in Appendix A.
@@
An allocation $f=\{f_{i}:i\in I\}$ is called a Walrasian equilibrium allocation for the exchange economy $\mathcal{E}$, if there exists a price vector $p\in\mathbb{R}^{d}$ such that $g_{i}\succ_{i}f_{i}$ implies that $p\,\cdot\,g_{i}>p\,\cdot\,\omega_{i}$, and $p\,\cdot\,f_{i}=p\,\cdot\,\omega_{i}$, for every $i\in I$.
An exchange economy $\mathcal{E}$ is convex if each preference relation $\succeq_{i}$ is convex, i.e., $\mathcal{E}\colon I\to\mathcal{C}\times\mathsf{E}^{d}_{+}$. Convexity is required for the basic theory of general equilibrium: existence of Walrasian equilibrium, as well as the second welfare theorem, relies on convex preferences. Our Theorems 2 and 3, will also rely on convexity of preferences.
When $s\mapsto\sum_{i\in I}\omega_{i}(s)$ is constant, we say that $\mathcal{E}$ exhibits no aggregate uncertainty. The aggregate endowment is then same across all states of the world, i.e., $\omega=(\bar{\omega},\ldots,\bar{\omega})$. The assumption of no aggregate uncertainty is common in the study of risk sharing: see for example Allen and Gale (1988) or, more closely related to our paper, Billot et al. (2000). Our Theorem 2 requires that there be no aggregate uncertainty.
@@ 3 Main Results
We proceed to state our main results, leaving a broader discussion of our findings to Section 4. The proofs of our results, including an overview of the methodology behind the proofs, can be found in Section 5.
@@ 3.1 Walrasian Equilibrium
Our first result is a generalization of the property that we have discussed in the introduction. It entails the following exercise: Fix a Walrasian equilibrium allocation, and draw a random perturbation from a ball of radius $r$. Consider the event that this random perturbation, when applied to any agent’s equilibrium consumption, results in contingent consumption that is better by at least $\varepsilon$ than their equilibrium consumption. The resulting event, that at least one agent is made better off, has vanishingly small probability.
@@
Let $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$ be an exchange economy. Suppose that there exists $\tau>0$ such that $\omega_{i}\geq\tau\bm{1}$ for all $i\in I$, and let $f$ be a Walrasian equilibrium allocation. Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Then for every $\varepsilon>0$,
@@
$\mathsf{P}^{r}\left((1-\varepsilon)(f_{i}+z)\succ_{i}f_{i}\text{ for some }i\in I\right)\leq\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}\,.$ (3.1)
@@
The proof of this theorem is in Section 5, as is the proof of the rest of the results we provide in this section. To unpack the statement of the theorem, imagine an economy with two consumers, $I=\{1,2\}$. If a random perturbation $z\in\mathbb{B}(r)$ is also in $(\mathcal{U}_{1}^{(\varepsilon)}(f_{1})-f_{1}\})\cup(\mathcal{U}_{2}^{(\varepsilon)}(f_{2})-\{f_{2}\})$, then it means that either $(1-\varepsilon)(z+f_{1})$ provides a strict welfare improvement for agent 1 over $f_{1}$, or that $(1-\varepsilon)(z+f_{2})$ does this for agent 2 over $f_{2}$ (or that both things are true). If we take $\tau=r$ and consider a 10% welfare improvement ($\varepsilon=0.1$), then the probability of making at least one agent better off is at most $e^{-d/800}$. Financial economists often take the state space to equal the returns to a set of assets that are traded in the market: Hence, if we take $d$ to be at least the number of stocks, say, trading on the NASDAQ Exchange, then our bound from Theorem 1 is about 1 %.
@@
We can express Theorem 1 in terms of volumes. Specifically, the bound in (3.1) is equivalent to
@@
By the monotonicity of preferences, $z>0$, a perturbation that leaves an agent with greater consumption state-by-state, is always a welfare improvement for any agent. Such a perturbation has probability $1/2^{d}$, which also decreases exponentially in the number of states. The “curvature” of agents’ preferences means, however, that many other perturbations should also result in a welfare improvement — recall the dotted area in Figure 1. So the conclusion in Theorem 1 may be surprising; it says that when the improvement is by a margin $\varepsilon$, the scope for welfare improvements shrinks exponentially in $d$ regardless of the shape of agents’ preferences, and even if we only ask for improvement in one agent’s welfare.
Finally, we should emphasize that Theorem 1 bounds the probability of an improvement to at least one agent, while the bound is independent of the number of agents. We may then consider an economy with many agents and many states, and conclude that it is very unlikely that a given perturbation could make even one agent better of by at least $\varepsilon$. Now if we do focus on a single agent, as we did in the introduction, then we obtain the statement that was discussed at length in the introduction:
@@
Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Under the hypotheses of Theorem 1, for every $\varepsilon>0$, and for each $i\in I$,
$\mathsf{P}^{r}\left((1-\varepsilon)(f_{i}+z)\succ_{i}f_{i}\right)\leq\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}\,.$ (3.3)
Equivalently, in terms of volumes, one has
We should emphasize that we interpret volume as a uniform probability measure on the ball of radius $r$, so that the statement in Corollary 1 corresponds to the informal description we provided in the introduction. The volume of the ball shrinks itself exponentially in $d$, so it is important to condition on the ball.
Observe that the message of Theorem 1 remains the same if we restrict attention to perturbations $z\in\mathbb{B}(f_{i},r)$ for which $p\cdot z>p\cdot\omega_{i}$. A consumption that is affordable cannot provide an improvement in utility over $f_{i}$, so it seems natural to consider only $z$ that costs more than $i$’s income at the equilibrium prices. This subset of $\mathbb{B}(f_{i},r)$ contains half its volume, so the message of Theorem 1 and Corollary 1 remains unchanged.
@@ 3.2 Optimality and resource utilization
Our second result considers collective welfare improvements, and concerns outcomes that may not be Walrasian equilibria. An equilibrium allocation must be Pareto optimal, but the Pareto set is strictly larger. So we ask about the possibility of a collective improvement after a change in aggregate consumption, starting from an allocation that is a weak Pareto optimum.
@@
Let $\mathcal{E}\colon I\to\mathcal{C}\times\mathbb{R}^{d}_{+}$ be a convex exchange economy, with no aggregate uncertainty, and aggregate endowment $\omega=\bm{1}$.
Assume $f$ is a weakly Pareto optimal allocation. For any given $\varepsilon>0$, let $\mathcal{V}^{(\varepsilon)}\coloneqq\sum_{i\in I}\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ be the Minkowski sum of the $\varepsilon$-approximate upper contour sets. Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Then
$\mathsf{P}^{r}\left(\sum_{i\in I}f_{i}+z\in\mathcal{V}^{(\varepsilon)}\right)\leq\mathrm{e}^{-\varepsilon^{2}d/8r^{2}}\,.$ (3.5)
@@
Similar to the Theorem 1, we can offer a version of the above result in the volume terms. Specifically, the bound in (3.5) is equivalent to
@@
$f(s_{1})$$f(s_{2})$$\omega$$f$$f(s_{1})$$f(s_{2})$$\omega$$f$
Figure 2: An illustration of Theorem 2
Figure 2 illustrates the concepts in the statement of Theorem 2. The figure depicts a case with two states, $s_{1}$ and $s_{2}$, and two individual traders. For pictorial simplicity we take $\varepsilon=0$ in the graphs. Aggregate consumption is $\omega=(1,1)$, and entails no aggregate uncertainty. The square defined by the origin and $\omega$ is an Edgeworth box that describes all the allocations between the two agents. The Pareto optimal allocations are on the solid curve that connects the origin and $\omega$, while $f=(f_{1},f_{2})$ is a particular Pareto optimal allocation (indicated by the tangency of the two agents’ indifference curves at the point $f$, with the usual Edgeworth-box convention that one agent’s coordinate system has been rotated $180^{\circ}$ and its origin coincides with $\omega$). For the allocation $f$ of the initial endowment $\omega$, we may consider all the aggregate consumptions that can be decentralized among the two agents so as to obtain higher utility than in $f$. This set of aggregate bundles, the Scitovsky contour $\mathcal{V}^{(0)}$, consists of all the bundles to the north-east of the dashed curve passing through $\omega$ (a curve that, at $\omega$, has the same slope as the common slope of each individual agent’s utility at $f$). [^6] The Scitovsky contour is a key concept in the proof of the second welfare theorem, where it plays a role that is analogous to our use in the proof of Theorem 2. For further discussion of Scitovsky contours and its connection to the Kaldor criterion and cost-benefit analysis see Debreu (1951), Samuelson (1956), or Echenique (2021).
We want to study the Scitovsky contour $\mathcal{V}^{(0)}$: understand and quantify how large it is. When agents’ utilities are monotone, $\mathcal{V}^{(0)}$ is, of course, an infinite set, and has infinite volume (Lebesgue measure; or area in the two-dimensional case). So we consider its volume relative to a ball centered at $\omega$. The exercise can be interpreted as follows: if we were to change aggregate consumption by randomly shifting $\omega$, what is the probability that we would end up in the Scitovsky contour. In other words, if we perturb $\omega$ by randomly (and uniformly) choosing a perturbation from the ball depicted in Figure 2, how likely is it that the resulting aggregate consumption would make both traders better off.
The exercise amounts to calculating the volume of the dotted region shown on the right of Figure 2. In the figure, the volume represents a significant fraction of the sphere centered in $\omega$. Less than $1/2$, but not by much. So it is quite likely that a random perturbation would make both agents better off. Theorem 2 says, however, that when $\varepsilon>0$ and the number of states is large, this volume will be negligible. It converges to zero very quickly as the number of states grows.
It is also worth emphasizing that the basic message of Theorem 2 holds, even if we restrict attention to perturbations that result in a more expensive aggregate bundle. For Figure 2, the probability of the dotted region once we condition on a more expensive perturbation is quite close to one. This probability, however, converges exponentially to zero in $d$. A similar point holds for Figure 1 in the Introduction.
@@
We shall reinterpret Theorem 2 in terms of Debreu’s coefficient of resource utilization (see Debreu (1951)). To this end, consider an allocation $f$ in an exchange economy $\mathcal{E}$ that is not weakly Pareto optimal. This means that there is an alternative allocation of the aggregate endowment in $\mathcal{E}$ that makes all agents strictly better off. In terms of the Scitovsky contour, this means that the aggregate endowment $\omega=\sum_{i}\omega_{i}$ lies in the set $\mathcal{V}^{(0)}$.
@@
Debreu considers a minimal amount of aggregate resources (call it $\omega^{*}$) that could be used to provide agents with the same utility as in $f$, and thinks of the gap between $\omega$ and $\omega^{*}$ as the inefficiency inherent in the allocation $f$. In Debreu’s words, these are “non-utilized resources.” He proposes to measure this gap by means of a “distance with economic meaning:” $p\cdot(\omega-\omega^{*})$, where $p$ is an “intrinsic price vector” associated with $\omega^{*}$. To obtain a measure that is, in a sense, scale independent, he works with the ratio of $p\cdot\omega^{*}$ to $p\cdot\omega$. Prices $p$ follow from an argument that is analogous to the second welfare theorem; in particular, they are not uniquely defined.
Debreu’s coefficient of resource utilization for an allocation $f=(f_{1},\ldots,f_{n})$ is defined as:
@@
$\mathrm{CRU}(f)\coloneqq\max_{\omega^{*}\in\partial\overline{\mathcal{V}^{(0)}}}\frac{p(\omega^{*})\cdot\omega^{*}}{p(\omega^{*})\cdot\omega},$
@@
Then we obtain, as a simple consequence of Theorem 2:
Corollary 2.
Let $\mathcal{E}\colon I\to\mathcal{C}\times\mathbb{R}^{d}_{+}$ be an exchange economy under the the hypotheses of Theorem 2.
Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. If $f$ is not weakly Pareto optimal, and $\mathrm{CRU}(f)$ its coefficient of resource utilization, then
$\mathsf{P}^{r}\left(\sum_{i\in I}f_{i}+z\in\mathcal{V}^{(0)}\right)\leq\mathrm{e}^{-(1-\mathrm{CRU}(f))^{2}d/8r^{2}}\,.$
Corollary 2 quantifies the meaning of the coefficient of resource utilization. Debreu writes that one may think of $\mathrm{CRU}(f)$ as a percentage of national income, or GDP. But in an economy with a large state space, even a seemingly large amount of inefficiency — as measured by the coefficient of resource utilization — may not translate into a wide scope for welfare improvements by changing aggregate consumption. To use the NASDAQ example from before, a seemingly large inefficiency of 50% measured by the $\mathrm{CRU}(f)$, translates into a Corollary 2 bound of $\mathrm{e}^{-112}$, which is essentially zero. In words, despite a large inefficiency of 50%, the chance that a random perturbation could be distributed to make all agents better off (not by some $\varepsilon>0$, just strictly better off) is essentially zero.

### v4

@@ 2.3 Exchange Economies
The set of agents is denoted by $I$ and a typical member of $I$ is referenced by index $i$. We assume throughout that $I$ is finite.
An exchange economy is a mapping $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$, where $\mathcal{E}(i)=(\succeq_{i},\omega_{i})$. Each agent $i\in I$ is described by a preference relation $\succeq_{i}$ on $\mathbb{R}^{d}$, as well as an endowment vector $\omega_{i}\in\mathbb{R}_{+}^{d}$. In an exchange economy, we use $\mathcal{U}_{i}^{(\varepsilon)}$ to denote the upper contour set $\mathcal{U}_{\succeq_{i}}^{(\varepsilon)}$.
Given an exchange economy $\mathcal{E}$, the aggregate endowment is $\omega\coloneq\sum_{i\in I}\omega_{i}$. A profile of acts across agents, say $f=(f_{i})_{i\in I}\in\mathbb{R}^{d\times I}$, is called an allocation if $\sum_{i\in I}f_{i}=\omega$. The space of all allocations is denoted by $\mathcal{F}_{\omega}$. We go back and forth between the textbook “allocation” terminology, and “risk-sharing agreement,” given our interpretation of consumption as a state-contingent monetary act. [^4] In the textbook Walrasian setting, $\omega(s)$ represents the total available amount of good $s$ in the economy, and in an Arrow-Debreu model it represents the total contingent amount that agents can receive in state $s$. See Chapter 7 in Debreu (1959), or Chapter 19.C in Mas-Colell et al. (1995).
@@
In words, an allocation $f$ is $\varepsilon$-Pareto optimal if there is no redistribution of the aggregate endowment $\omega=\sum_{i}f_{i}$ that would be strictly better for all agents, and that would remain strictly better for all agents after a fraction $\varepsilon$ of consumption is “shaved off” in each state of the world. When $\varepsilon=0$, the definition coincides with the usual notion of weak Pareto optimality. When we insist on the risk-sharing interpretation of our model, Pareto optimality means an ex-ante efficient risk-sharing agreement.
@@
An allocation $f=(f_{i})_{i\in I}$ is called a Walrasian equilibrium allocation for the exchange economy $\mathcal{E}$, if there exists a price vector $p\in\mathbb{R}^{d}$ such that $g_{i}\succ_{i}f_{i}$ implies that $p\,\cdot\,g_{i}>p\,\cdot\,\omega_{i}$, and $p\,\cdot\,f_{i}=p\,\cdot\,\omega_{i}$, for every $i\in I$.
An exchange economy $\mathcal{E}$ is convex if each preference relation $\succeq_{i}$ is convex, i.e., $\mathcal{E}\colon I\to\mathcal{C}\times\mathbb{R}_{+}^{d}$. Convexity is required for the basic theory of general equilibrium: existence of Walrasian equilibrium, as well as the second welfare theorem, relies on convex preferences. Our Theorems 2 and 3 will also rely on the convexity of agents’ preferences (Theorem 1 however, makes no such assumption, but it is a statement about Walrasian equilibrium allocations).
When $s\mapsto\sum_{i\in I}\omega_{i}(s)$ is constant, we say that $\mathcal{E}$ exhibits no aggregate uncertainty. The aggregate endowment is then the same across all states of the world, i.e., $\omega=(\bar{\omega},\ldots,\bar{\omega})$. The assumption of no aggregate uncertainty is common in the study of risk sharing: see, for example, Mas-Colell, Whinston, and Green (1995, Ch. 19.C)Allen and Gale (1988) or, more closely related to our paper, Billot, Chateauneuf, Gilboa, and Tallon (2000). Our Theorem 2 requires that there be no aggregate uncertainty.
@@ 3 Main Results
We proceed to state our main results. A broader discussion of our findings can be found in the introduction. We discuss more specific questions of interpretation in Section 4. The proofs of our results, including an overview of the methodology behind the proofs, can be found in Section 5.
@@ 3.1 Walrasian Equilibrium
Our first result entails the following exercise: Fix a Walrasian equilibrium allocation, and draw a random perturbation from a ball of radius $r$ distributed according to some probability measure $\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. [^5] Obviously the interesting measures $\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$ involve $\kappa$ that is independent of $d$, which is the case for the uniform and conditional Gaussian measures. See our remarks in Section 2.1. Consider the event that this random perturbation, when applied to any agent’s equilibrium consumption, results in a contingent consumption that is better by at least $\varepsilon$ than their equilibrium consumption. The resulting event, that at least one agent is made better off, has vanishingly small probability.
@@
Let $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$ be an exchange economy. Suppose that there exists $\tau>0$ such that $\omega_{i}\geq\tau\bm{1}$ for all $i\in I$, and let $f$ be a Walrasian equilibrium allocation. Fix $r>0$ and let $z\sim\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. Then for every $\varepsilon>0$,
@@
$\mathsf{P}^{r}_{\kappa}\left((1-\varepsilon)(f_{i}+z)\succ_{i}f_{i}\text{ for some }i\in I\right)\leq\kappa\,\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}\,.$ (3.1)
The proof of this theorem is in Section 5, as is the proof of the rest of the results we provide in this section. The proof relies on results from the literature on isoperimetric inequalities and the concentration of measure phenomenon, but the proof section attempts to be self-contained. [^6] One simple intuition highlights the connection with the concentration of measures: Consider an economy with a single risk-neutral agent, and assume that all states are equally likely; so the agent’s preferences are represented by $u(f)=\sum_{s}f(s)$. Then an $\varepsilon$-improvement over a constant act $f(s)=1$ (which would be an equilibrium allocation in a Robinson Crusoe economy) requires that $(1-\varepsilon)\sum_{s}z(s)>\varepsilon d$, or that $\frac{1}{d}\sum_{s}z(s)>\varepsilon/(1-\varepsilon)$. In high dimensions, despite the lack of independence ($z$ is drawn from a sphere), the random variable $\frac{1}{d}\sum_{s}z(s)$ is tightly concentrated around 00.
$f(s_{1})$$f(s_{2})$$f_{i}$$\omega_{i}$$p$
$f(s_{1})$$f(s_{2})$$f_{i}$$\omega_{i}$$p$
Figure 2: Individual welfare.
To unpack the statement of Theorem 1, imagine an economy with two consumers, $I=\{1,2\}$. The situation of an individual agent $i$ is illustrated by Figure 2, which shows the usual consumption-maximization diagram. In the figure, there are two states of the world ($d=2$), and an agent’s budget set is depicted in gray. The budget is defined by an equilibrium price $p$, and by the agent’s endowment $\omega_{i}$. The picture shows the agent’s equilibrium consumption $f_{i}$. The agent’s indifference curve, tangent to the budget line at $f_{i}$, is depicted as a dashed curve. The upper contour set is the region to the northeast of the indifference curve that contains all consumptions that are preferred to $f_{i}$. If a random perturbation $z\in\mathbb{B}(r)$ is in $(\mathcal{U}_{1}^{(\varepsilon)}(f_{1})-f_{1}\})\cup(\mathcal{U}_{2}^{(\varepsilon)}(f_{2})-\{f_{2}\})$, then it means that either $(1-\varepsilon)(z+f_{1})$ provides a strict welfare improvement for agent 1 over $f_{1}$, or that $(1-\varepsilon)(z+f_{2})$ does this for agent 2 over $f_{2}$ (or that both things are true).
The consumer would be better off with more consumption in every state, but they would also be willing to accept a tradeoff of more consumption in one state than in another. Of course, the tradeoffs would have to be at terms of trade that are more favorable than equilibrium prices, and guided by the curvature of the agent’s utility function (and its implied level of risk aversion). But the figure shows that there is a significant probability that a perturbation would leave the agent strictly better off, and the upper contour set contains much more than the bundles that are larger state-by-state. So does the dotted region. The theorem, actually, deals with the event that $z$ is in $(\mathcal{U}_{1}^{(\varepsilon)}(f_{1})-f_{1}\})\cup(\mathcal{U}_{2}^{(\varepsilon)}(f_{2})-\{f_{2}\})$, a potentially non-convex set that is usually strictly larger than the dotted region (and even larger when there are more than two agents).
The implications of the theorem may, therefore, be surprising for reasons that are familiar from our discussion in the introduction. The perturbations $z$ with $(1-\varepsilon)(f_{i}+z)\succ_{i}f_{i}$ do not all feature $z>0$. In fact, the figure, as well as basic economic intuition, suggests that the curvature of the agents’ utility functions should matter in calculating the probability in Theorem 1. So it is surprising that, for large $d$, the curvature, or risk attitude, does not matter and we get an exponentially decreasing probability of a welfare improvement; the same as we would get if we were dealing with extremely risk-averse agents, and asking for a perturbation that is positive in all dimensions.
@@
To put the bound in the theorem in context, consider an example with $\tau=r$, a uniform probability measure, and a 11% welfare improvement ($\varepsilon=0.1$). Then the probability of making at least one agent better off is at most $e^{-d/800}$. Financial economists often take the state space to equal the returns to a set of assets that are traded in the market: Hence, if we take $d$ to be the number of stocks, say, trading on the NASDAQ Exchange, then our bound from Theorem 1 is about 0.67%. [^7] The number of stocks listed on NASDAQ is, according to https://www.nasdaq.com/, at least 4000.
@@
We can express Theorem 1 in terms of volumes. Specifically, when $z$ is drawn from the uniform measure $\mathsf{P}^{r}$, the bound in (3.1) is equivalent to
@@
Finally, we should emphasize that Theorem 1 bounds the probability of an improvement to at least one agent, while the bound is independent of the number of agents. We may then consider an economy with many agents and many states, and conclude that it is very unlikely that a given perturbation could make even one agent better off by at least $\varepsilon$.
Now, if we do focus on a particular agent, we immediately obtain:
@@
Fix $r>0$ and let $z\sim\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. Under the hypotheses of Theorem 1, for every $\varepsilon>0$, and for each $i\in I$,
$\mathsf{P}^{r}_{\kappa}\left((1-\varepsilon)(f_{i}+z)\succ_{i}f_{i}\right)\leq\kappa\,\mathrm{e}^{-\varepsilon^{2}\tau^{2}d/8r^{2}}\,.$ (3.3)
Observe that the message of Theorem 1 remains the same if we restrict attention to perturbations $z\in\mathbb{B}(f_{i},r)$ for which $p\cdot z>p\cdot\omega_{i}$. In a Walrasian equilibrium, the set $\bigcup_{i\in I}\left(\mathcal{U}_{i}^{(\varepsilon)}(f_{i})-f_{i}\right)$ lies in the half-space $\{z:p\cdot z\geq 0\}$, as $p\cdot f_{i}=p\cdot\omega_{i}$, and a consumption that is affordable cannot provide an improvement in utility over $f_{i}$. So it seems natural to consider only $f_{i}+z$ that costs more than $i$’s income at the equilibrium prices. This subset of $\mathbb{B}(r)$ contains half its volume, and thus the message of Theorem 1 and Corollary 1 remains unchanged.
@@ 3.2 Optimality and Resource Utilization
Our second result considers collective welfare improvements, and concerns outcomes that may not be Walrasian equilibria. An equilibrium allocation must be Pareto optimal, but the Pareto set is strictly larger. We ask about the possibility of a collective improvement after a shock in aggregate consumption, starting from an allocation that is a weak Pareto optimum.
@@
Let $\mathcal{E}\colon I\to\mathcal{C}\times\mathbb{R}_{+}^{d}$ be a convex exchange economy with no aggregate uncertainty, and an aggregate endowment of $\omega=\bm{1}$. Assume that $f$ is a weakly Pareto optimal allocation in $\mathcal{E}$. For any $\varepsilon>0$, let $\mathcal{V}^{(\varepsilon)}(f)\coloneq\sum_{i\in I}\mathcal{U}_{i}^{(\varepsilon)}(f_{i})$ be the Minkowski sum of the $\varepsilon$-upper contour sets. Fix $r>0$ and let $z\sim\mathsf{P}^{r}_{\kappa}\in\mathcal{M}^{r}_{\kappa}$. Then
$\mathsf{P}^{r}_{\kappa}\left(\omega+z\in\mathcal{V}^{(\varepsilon)}(f)\right)\leq\kappa\,\mathrm{e}^{-\varepsilon^{2}d/8r^{2}}\,.$ (3.4)
The set $\mathcal{V}^{(0)}(f)$ of all aggregate bundles that may be disaggregated to make all agents strictly better off is called the Scitovsky contour; it is a basic notion in welfare economics (see De Scitovszky (1942) and Samuelson (1956)), and plays a key role in the proof of the second welfare theorem. Here we focus on a version that requires $\varepsilon$-welfare improvements.
@@
Similar to Theorem 1, we can offer a version of the above result in terms of volume. Specifically, under the uniform measure $\mathsf{P}^{r}$, the bound in (3.4) is equivalent to
@@
It is also worth emphasizing that, just like in Theorem 1, the basic message of Theorem 2 holds, even if we restrict attention to perturbations that result in a more expensive aggregate bundle. So, if $p$ is the supporting price vector obtained from the second welfare theorem, we can restrict to perturbations $z$ with $p\cdot z>0$. These are the perturbations that, according to neoclassical cost-benefit analysis (or the Kaldor criterion, see Graaff (1967)), have some chance of being Pareto improving. The reason is the same as that previously discussed for Theorem 1.
@@
A version of Theorem 2 holds for allocations that are not Pareto optimal. We make use of the measure of inefficiency introduced by Debreu (1951), the coefficient of resource utilization. To this end, consider an allocation $f$ in an exchange economy $\mathcal{E}$ that is not weakly Pareto optimal. This means that there is an alternative allocation of the aggregate endowment in $\mathcal{E}$ that makes all agents strictly better off. In terms of the Scitovsky contour, this means that the aggregate endowment $\omega=\sum_{i}\omega_{i}$ lies in the set $\mathcal{V}^{(0)}$.
@@
Debreu considers a minimal amount of aggregate resources (call it $\omega^{*}$) that could be used to provide agents with the same utility as in $f$, and thinks of the gap between $\omega$ and $\omega^{*}$ as the inefficiency inherent in the allocation $f$. In Debreu’s words, these are the “non-utilized resources.” He proposes to measure this gap by means of a “distance with economic meaning:” $p\cdot(\omega-\omega^{*})$, where $p$ is an “intrinsic price vector” associated with $\omega^{*}$. To obtain a measure that is, in a sense, scale independent, he actually works with the ratio of $p\cdot\omega^{*}$ to $p\cdot\omega$. Prices $p$ follow from an argument that is analogous to the second welfare theorem; in particular, they are not uniquely defined.
Debreu’s coefficient of resource utilization for an allocation $f=(f_{i})_{i\in I}$ is defined as:
@@
$\mathrm{CRU}(f)\coloneq\max_{\omega^{*}\in\partial\overline{\mathcal{V}^{(0)}}}\frac{p(\omega^{*})\cdot\omega^{*}}{p(\omega^{*})\cdot\omega},$
@@
Then we obtain, following a similar argument as in Theorem 2:
Proposition 1.
Let $\mathcal{E}\colon I\to\mathcal{C}\times\mathbb{R}_{+}^{d}$ be an exchange economy under the hypotheses of Theorem 2. Fix $r>0$ and let $z\sim\mathsf{P}^{r}_{\kappa}$. If $f$ is not weakly Pareto optimal, and $\beta\coloneq\mathrm{CRU}(f)$ is its coefficient of resource utilization, then
$\mathsf{P}^{r}_{\kappa}\left(\omega+z\in\mathcal{V}^{(1-\beta^{2})}(f)\right)\leq\mathrm{e}^{-\left(\frac{1-\beta}{\beta}\right)^{2}d/8r^{2}}\,.$ (3.6)
Proposition 1 quantifies the meaning of the coefficient of resource utilization. Debreu writes that one may think of $\mathrm{CRU}(f)$ as a percentage of national income, or GDP. But in an economy with a large state space, even a seemingly large amount of inefficiency—as measured by the coefficient of resource utilization—may not translate into a wide scope for welfare improvements by changing aggregate consumption. To use a similar numerical example from before, consider the case where $d$ is calculated from the number of stocks listed on NASDAQ. Suppose an inefficiency of 10%, meaning a CRU of 0.9. Then the welfare improvement in Proposition 1 is about 19% and the bound on the probability of such an improvement is about 0.21%.